OpenStudy (itsbribro):

which equation is in point-slope form for the given point and slope? point: (5, 9) slope: 2

OpenStudy (itsbribro):

A. y-9=2(x+5) B. y-5=2(x+9) C. y-9=2(x-5) D. y+9=2(x-5)

OpenStudy (itsbribro):

@ninjaslice @carlyleukhardt @AloneS

OpenStudy (ninjaslice):

can you please show the point and the slope?

OpenStudy (itsbribro):

point: (5, 9) and the slope is 2

OpenStudy (itsbribro):

.... please help :'(